OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x^4 + 49x^2

OpenStudy (harsimran_hs4):

straight away i see x^2 can be taken out as common to write x^2(x^2 + 49) x^2 + 49 will have complex roots if you are working in real plane factorization is (x)(x)(x^2 + 49)

OpenStudy (anonymous):

then would you make it into x^4+49?

OpenStudy (harsimran_hs4):

@Andreshoyos23 you mean to say the final answer would be x^4+49 ?

OpenStudy (anonymous):

yes ? im a bit confused with what im supposed to do .

OpenStudy (anonymous):

@harsimran_hs4

OpenStudy (harsimran_hs4):

no lets check it step by step = x^4 + 49x^2 = x(x^3 + 49x) = x^2(x^2 + 49) clear till now ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@harsimran_hs4 but then where would u go from here? i need to work it into an imaginary number but i dont know how to do that would it just factor the x^2 and 49? or foil i forget the names.

OpenStudy (harsimran_hs4):

ok now we had reached x^2(x^2 + 49) now forget every thing and tell me the values of x which satisfy x^2 + 49 = 0

OpenStudy (anonymous):

@harsimran_hs4 (+- 7)?

OpenStudy (anonymous):

@harsimran_hs4 7i

OpenStudy (harsimran_hs4):

yes +7i and -7i so finally x^2 (x^2 + 49) = x^2(x - 7i) (x+7i) is the final answer

OpenStudy (anonymous):

the one - is because when u place it on the other side it becomes negative and you have to show that in the equation right ? and thank you so much u were lots of help